DETERMINATION OF THE POTENTIAL ENERGY CURVE OF THE D^{1}Π STATE OF NaK BY TWO-COLOR POLARIZATION LABELING SPECTROSCOPY

\underline{AMANDA ROSS}, Université Lyon 1 and CNRS (LASIM, UMR 5579), 69622 Villeurbanne, France; ADJO ADOHI-KROU, UFR SSMT Université de Cocody, 22 BP 582 Abidjan, Ivory Coast Republic; WLODZIMIERZ JASTRZȨBSKI, Institute of Physics, Polish Academy of Sciences, Al. Lotnikóv 32/46, 02-668 Warsaw, Poland; PAWEŁ KOWALCZYK, Institute of Experimental Physics, Warsaw University, ul. Hoża 69, 00-681 Warsaw, Poland; and ANDREY V. STOLYAROV, Department of Chemistry, Moscow State University, 119992 Moscow, Russia.

Two-colour polarization labeling experiments measuring the D ← X system of NaK have furnished observations of the D ^{1}Π state of NaK up to v” = 42. The last observed level is located 7 cm^{−1} below the Na(3p ^{2}P_{3/2}) + K(4s) atomic asymptote, clearly indicating the dissociation products of this state. The vibrational progressions all exhibit irregular intervals, predominantly because of strong interactions with the nearby d ^{3}Π state, which also dissociates to Na(3p) + K(4s) atoms. The polarization data have been combined with some resolved fluorescence D → X transitions, and analyzed by fitting to spectroscopic parameters and to an analytical potential curve. A full deperturbation treatment has not been attempted, but a “robust” weighting scheme has been used to reduce the influence of levels that cannot be properly represented by a single channel model. Parameters determined in a fit to a potential curve include T_{e} = 20090.15 ± 0.02 cm^{−1}, D_{e} = 2157.0 ± 0.3 cm^{−1}, R_{e} = 4.155 ± 0.002 Å, with an unweighted root mean square error of 0.12 cm^{−1} for 959 data.